Solve the following equation & list all possible solutions

6cos^2x-13sinx-13=0
Answer #1

Solution a 6 cos? cge) -13 Sin (s)-13=0 6 (I sin²(x)) -13 sinc&)-13 = 0 6- 6 sin²(&) -13 sin (2) +13=0 6 sin?().+ 13 sin(x) +Please Upvote this